SC reports the following income statement results:

Sales

$907,000

Operating expense

288,400

Net income

94,500

Sales returns & allowances

7,000

Gross profit

401,400

Interest revenue

4,000

  1. $____________Calculate Net sales:                                                               

  1. $____________Calculate Cost of Goods Sold                                             

  1. $____________Calculate operating income                                                

  1. $_____________Calculate Income before Income tax (IBT)                

  1. ___ ___.___%.  Calculate the gross profit margin(one decimal place)

__ ___.___%.  Calculate the net profit margin (one decimal place
Answer #1

Answer:

Net sales $    900,000.00
Less: Cost of Goods Sold $    498,600.00
Operating Income $    113,000.00
Income before taxes IBT $    109,000.00

Gross Margin = $ 401,400 / 900,000

= 44.60%

Net Profit Margin Ratio = $ 94,500 / 900,000

= 10.50%

Working Note:

Gross sales (A) $    907,000.00
Less: Sales returns (B) $        7,000.00
Net sales (A)-(B)=C $    900,000.00
Less: Cost of Goods Sold (D-C) $    498,600.00
Gross Profit (D) $    401,400.00
Less: Operating expenses (E $    288,400.00
EBIT (D-E)=(F) $    113,000.00
Less: Interest (G) $        4,000.00
Income before taxes IBT (F-G)=(H) $    109,000.00
Less: taxes (H-I)=(J) $      14,500.00
Net Income (I) $      94,500.00

In case of any doubt, please comment.